Minutes of the Trigger Operation Meeting (June 27, 2006)
Attendees: Al Eisner, Fang Fang, Selina Li, Rahmat Rahmat, Dong Su
- Smooth data-taking except some downtime from PEP over this weekend and yesterday.
- Yesterday, the shifter noticed that the EMT phi vs. theta plot had a noisy bin in run 66096 but it seems to correspond to the EMC hot channels in that region.
- DCZ failed to configure after PEP's timing signal was lost on June 18. Selina noticed that one of the boards in TSF-Y wasn't detected at configuration. Power cycling the DCZ crates solved the problem.
- There was a note in the elog about timing fiducial errors, but no mention of problem with the DCZ. Trigger data taken after the timing fiducial problem was solved, so we assume TRG didn't get affected.
Data Quality and Monitoring
- Rahmat continues to work on the lumi ratio.
- Selina showed some plots obtained from running Rainer's trigger statistics tool for the pi+ pi- Run5 compared to Run4, pi0 pi0, p+ p- signal MC samples in two-photon physics. The comparison of the run4 and run5 plots seems to indicate a ~2% lost in 1Z' efficiency, so Su Dong suggested to look at these plots separately for different W regions to see where we lose these events.
This page is created by Selina Li (firstname.lastname@example.org); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.